OpenStudy (anonymous):

A poster of total area 256 in2 is to have a margin of 4 inches at the top and bottom and 1 inch at each side. (a) Find the dimensions of the poster which give the largest printed area. I got: 32 height in; 8 width in But I'm having trouble with this part. (b) Suppose that, instead, we know the poster will contain 256 in2 of printed material with margins of 4 inches at the top and bottom and 1 inch at each side. Find the dimensions of the poster that minimize its total size. height in width in

OpenStudy (anonymous):

But I got the first part right, so should I treat each margin as it's own thing like i did in the other one?

OpenStudy (anonymous):

you know that (A) is correct with your answers? If so, I may not have the correct approach.

OpenStudy (anonymous):

Yes I got green check marks lol. It's an online math hw site

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

@byerskm2 For part B, try it like this... Let the height and width of the printed part be h and w. So h * w = 256 sq in. Then, the outside area of the whole poster is A = (h+8)(w+2) Then substitute in using h = 256/w from the first equation. This will allow you to get an equation of function for the outside area all in the variable w... A(w), in other words. Then find dA(w)/dw, and solve for w that makes the derivative equal 0. Then go back and solve for h.